HELPMEASAPPP:

If log 2=0.3010 and log 3=0.4771, find the value of: 1) log 6 2) log 5

LKylie:

i) log 6= log 2 x 3 = log 2 + log 3 = 0.3010+0.4771 = 0.7781 ii) log 5= log 10/2 = log 10 - log 2 = 1-0.3010= 0 6990

HELPMEASAPPP:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @LKylie i) log 6= log 2 x 3 = log 2 + log 3 = 0.3010+0.4771 = 0.7781 ii) log 5= log 10/2 = log 10 - log 2 = 1-0.3010= 0 6990 \(\color{#0cbb34}{\text{End of Quote}}\) Thnx

LKylie:

Np